# Optimal Transport Guided Correlation Assignment for Multimodal Entity Linking

### v1

@@ Abstract
Zefeng Zhangthanks:  These authors contribute equally to this work., Jiawei Sheng11footnotemark: 1, Chuang Zhangthanks:  Corresponding Authors., Yunzhi Liang, Wenyuan Zhang, Siqi Wang, Tingwen Liu22footnotemark: 2 Institute of Information Engineering, Chinese Academy of Sciences. Beijing, China School of Cyber Security, University of Chinese Academy of Sciences. Beijing, China {zhangzefeng,shengjiawei,zhangchuang}@iie.ac.cn {liangyunzhi,zhangwenyuan,wangsiqi2022,liutingwen}@iie.ac.cn
@@ 1 Introduction
Entity Linking (EL) is an important yet challenging task in knowledge acquisition, which can facilitate applications such as information retrieval Liu et al. (2018), question answering Xiong et al. (2019) and dialogue systems Jiang et al. (2023). In recent years, with the rapid increase of multimodal content, Multimodal Entity Linking (MEL) has been proposed, which aims to link ambiguous mentions in multimodal contexts to entities in the multimodal knowledge graph (KG) (Sheng et al., 2020; Li et al., 2023). As shown in Figure 1, given a mention Robert in texts with related images, the MEL model requires the mention to be corresponded to the entity Robert_Downey in the multimodal KG.
@@ 3.1 Multimodal Feature Extraction
We employ the pre-trained BERT to extract textual features. For mention $m$, we concatenate the word of mention $m_{n}$ with its context sentence $m_{c}$, and feed it into BERT as:
@@ 5 Related Works
Traditional EL mainly focuses on the textual information of mentions and entities. These methods Cao et al. (2017); Wu et al. (2019) first employ textual encoders to obtain contextual representations of mention and entity, and then calculate similarities between them as final probabilities. To enhance consistency in EL, they enrich local correlations Peters et al. (2019); Liu et al. (2023) with prior knowledge and collective mechanism, and global correlations Le and Titov (2018); Cao et al. (2018); Fang et al. (2019) with topical coherence among mentions. However, these methods are designed to deal with text only, which can hardly deal with the crucial multimodal correlations within images in MEL. Other studies have focused on Generative Entity Linking(GEL) (De Cao et al., 2020; Shi et al., 2023; De Cao et al., 2022; Huang et al., 2022; Yuan et al., 2022). This is a promising research problem in the era of Large Language Models (LLMs), but they suffer from lexical ambiguity problem (Fortuny and Payrató, 2024), which we discuss in detail in the Appendix G.
@@
Recent years have seen numerous social media and news posted with texts and images, thereby MEL is proposed to identify interested entities from multimodal information. As a pioneering study, Moon et al. (2018) firstly introduces images to benefit EL, which blends visual, word and character features through an attention mechanism. Afterward, there exist several studies Adjali et al. (2020); Zheng et al. (2022); Wang et al. (2022a); Luo et al. (2023) exploring gate or attention mechanism to capture complicated multimodal correlations for mentions and entities. Recently, Xing et al. (2023) introduce external object-level scene graphs of images to enrich visual correlations, but may involve elaborate errors and still construct multimodal correlations with attentive graph neural networks. Luo et al. (2023) present a sophisticated gated framework for fine-grained inter-modal interactions, achieving the most advanced results. Although existing studies attempt multiple correlations in different aspects, they heavily rely on the automatically learned attention-based mechanism, which can hardly ensure appropriate correlation assignments and impede further improvements.
@@ 6 Conclusion
In this paper, we propose a MEL model with OT correlation assignment, namely OT-MEL. First, OT-MEL simplistically assumes uniform distributions for both source and target, which can be further explored in real-world scenarios. Additionally, as OT-MEL is tailored for visual and textual modality in MEL, it doesn’t consider OT in other modalities such as speech and video. We will explore the generality of the proposed OT correlation assignment in these complex modalities for wide research communities.
@@ Appendix D Metrics
We first calculate the similarity scores between a mention and all entities of the KB, then the similarity scores are sorted in descending order to calculate H@k and MRR, which are defined as:
@@ Appendix G Dissusion of the Generative Entity Linking
With the emergence of LLMs, the GEL is a promising paradigms nowadays. However, how to solve the lexical ambiguity problem of GEL is one of the most important problems in its application, and we will discuss it in the future work.

### v2

@@ Abstract
Zefeng Zhangthanks:  These authors contribute equally to this work., Jiawei Sheng11footnotemark: 1, Chuang Zhangthanks:  Corresponding Authors., Yunzhi Liang, Wenyuan Zhang, Siqi Wang, Tingwen Liu Institute of Information Engineering, Chinese Academy of Sciences. Beijing, China School of Cyber Security, University of Chinese Academy of Sciences. Beijing, China {zhangzefeng,shengjiawei,zhangchuang}@iie.ac.cn {liangyunzhi,zhangwenyuan,wangsiqi2022,liutingwen}@iie.ac.cn
@@ 1 Introduction
Entity Linking (EL) is an important yet challenging task in knowledge acquisition, which can facilitate applications such as information retrieval Liu et al. (2018), question answering Xiong et al. (2019) and dialogue systems Jiang et al. (2023). In recent years, with the rapid increase of multimodal content, Multimodal Entity Linking (MEL) has been proposed, which aims to link ambiguous mentions in multimodal contexts to entities in the multimodal knowledge graph (KG). As shown in Figure 1, given a mention Robert in texts with related images, the MEL model requires the mention to be corresponded to the entity Robert_Downey in the multimodal KG.
@@ 3.1 Multimodal Feature Extraction
We employ BERT to extract textual features. For mention $m$, we concatenate the word of mention $m_{n}$ with its context sentence $m_{c}$, and then feed it into BERT as:
@@ 5 Related Works
Traditional EL mainly focuses on the textual information of mentions and entities. These methods Cao et al. (2017); Wu et al. (2019) first employ textual encoders to obtain contextual representations of mention and entity, and then calculate similarities between them as final probabilities. To enhance consistency in EL, they enrich local correlations Peters et al. (2019); Liu et al. (2023) with prior knowledge and collective mechanism, and global correlations Le and Titov (2018); Cao et al. (2018); Fang et al. (2019) with topical coherence among mentions. However, these methods are designed to deal with text only, which can hardly deal with the crucial multimodal correlations within images in MEL. Other studies have focused on Generative Entity Linking (GEL) (De Cao et al., 2020; Shi et al., 2023; De Cao et al., 2022; Huang et al., 2022; Yuan et al., 2022). This is a promising research problem in the era of Large Language Models (LLMs), but they can suffer from the lexical ambiguity problem (Fortuny and Payrató, 2024), which we will discuss in details in the Appendix G.
@@
Recent years have seen numerous social media and news posted with texts and images, thereby MEL is proposed to identify interested entities from multimodal information, which can benefit to knowledge acquisition (Sheng et al., 2020; Li et al., 2023). As a pioneering study, Moon et al. (2018) firstly introduces images to benefit EL, which blends visual, word and character features through an attention mechanism. Afterward, there exist several studies Adjali et al. (2020); Zheng et al. (2022); Wang et al. (2022a); Luo et al. (2023) exploring gate or attention mechanism to capture complicated multimodal correlations for mentions and entities. Recently, Xing et al. (2023) introduce external object-level scene graphs of images to enrich visual correlations, but may involve elaborate errors and still construct multimodal correlations with attentive graph neural networks. Luo et al. (2023) present a sophisticated gated framework for fine-grained inter-modal interactions, achieving the most advanced results. Although existing studies attempt multiple correlations in different aspects, they heavily rely on the automatically learned attention-based mechanism, which can hardly ensure appropriate correlation assignments and impede further improvements.
@@ 6 Conclusion
In this paper, we propose a MEL model with OT correlation assignment, namely OT-MEL. First, OT-MEL simplistically assumes uniform distributions for both source and target, which can be further explored in real-world scenarios. Besides, as OT-MEL is tailored for visual and textual modality in MEL, it doesn’t consider OT in other modalities such as speech and video. We will explore the generality of the proposed OT correlation assignment in these complex modalities for wide research communities.
Acknowledgement
We would like to thank the anonymous reviewers for their comments. This work was supported by the National Key Research and Development Program of China (Grant No.2021YFB3100600), the Youth Innovation Promotion Association of CAS (No.2021153), and the Postdoctoral Fellowship Program of CPSF (No.GZC20232968).
@@ Appendix D Metrics
We first calculate the similarity scores between a mention and all entities of the KG, then the similarity scores are sorted in descending order to calculate H@k and MRR, which are defined as:
@@ Appendix G Dissusion of the Generative Entity Linking
With the emergence of LLMs, the GEL is a promising paradigms nowadays. However, how to solve the lexical ambiguity problem of GEL is one of the most important problems in its application, and we will explore it in the future work.